### v4

@@ B.1   Non-binary, ordered treatment
For the numerator in the Wald DID estimand, we have
@@
The first equality follows from the simple manipulation and Assumption 4. The second equality follows from Assumption 6.
@@
For the denominator in the Wald-DID estimand, by similar argument, we obtain:
@@
The first equality follows from simple algebra and Assumption 4. The second equality follows from Assumption 5.
@@
B.2   Repeated cross sections
Here, we present our identification results under repeated cross section settings.
B.2.1 Two time periods
Let $Y_{i}$ and $D_{i}$ denote the outcome and the treatment for unit $i$. Let $Z_{i}$ denote the instrument path for unit $i$: $Z_{i}=(0,0)$ if unit $i$ is not exposed to the instrument and $Z_{i}=(0,0)$ if unit $i$ is exposed to the instrument. Let $E_{i}\in\{0,1\}$ denote the group indicator for unit $i$: $E_{i}=1$ if $Z_{i}=(0,1)$ and $E_{i}=0$ if $Z_{i}=(0,0)$. Let $T_{i}\in\{0,1\}$ denote the binary indicator for time. For all $z\in\mathcal{S}(Z)$, let $Y_{i}(0,z),Y_{i}(1,z)$, and $D_{i}(z)$ denote potential outcomes and potential treatment choices for unit $i$. Let $D_{i,t}(z)$ denote potential treatment choices under $T_{i}=t$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
In a two-period and two-group setting, our target parameter is the local average treatment effect on the treated in period $1$:
@@ B.2.1 Two time periods
Theorem 5 below shows that the Wald-DID estimand is equal to the LATET in period $1$ under Assumptions 14-18. Here, we assume the binary treatment, but the non-binary, ordered treatment case is easy to extend as in Theorem 3 (thus, we omit it for brevity). In the proof, we omit the index $i$ to ease the notation.
Theorem 5.
Suppose Assumptions 14-18 hold. If we assume a binary treatment, the Wald-DID estimand corresponds to the local average treatment effect on the treated (LATET) in period $1$:
$w_{DID}$ $\equiv\frac{E[Y|E=1,T=1]-E[Y|E=1,T=0]-\{E[Y|E=0,T=1]-E[Y|E=0,T=0]\}}{E[D|E=1,T=1]-E[D|E=1,T=0]-\{E[D|E=0,T=1]-E[D|E=0,T=0]\}}$
$=E[Y(1)-Y(0)|T=1,E=1,CM^{Z}]$
$=LATET.$
Proof.
First we consider the denominator of the Wald-DID estimand. For the denominator, we have
$E[D|E=1,T=1]-E[D|E=1,T=0]-\{E[D|E=0,T=1]-E[D|E=0,T=0]\}$
$=$ $E[D((0,1))|E=1,T=1]-E[D((0,1))|E=1,T=0]$
$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}$
$=$ $E[D((0,1))-D((0,0))|E=1,T=1]$
+$+\Bigg{\{}E[D((0,0))|E=1,T=1]-E[D((0,1))|E=1,T=0]$
$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}\Bigg{\}}$
$=$ $E[D((0,1))-D((0,0))|E=1,T=1]$
+$+\Bigg{\{}E[D((0,0))|E=1,T=1]-E[D((0,0))|E=1,T=0]$
$-\{E[D((0,0))|E=0,T=1]-E[D((0,0))|E=0,T=0]\}\Bigg{\}}$
$=$ $Pr(D_{1}((0,1))>D_{1}((0,0))|E=1,T=1)$
$=$ $Pr(CM^{Z}|E=1,T=1).$ (14)
The second equality follows from the simple manipulation. The third equality follows from Assumption 16 and the fourth equality follows from Assumption 15 and Assumption 17.
Next, for the numerator, we have
$E[Y|E=1,T=1]-E[Y|E=1,T=0]-\{E[Y|E=0,T=1]-E[Y|E=0,T=0]\}$
$=$ $E[Y(D((0,1)))|E=1,T=1]-E[Y(D((0,1)))|E=1,T=0]$
$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\}$
$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
+$+\Bigg{\{}E[Y(D((0,0)))|E=1,T=1]-E[Y(D((0,1)))|E=1,T=0]$
$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\Bigg{\}}$
$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
+$+\Bigg{\{}E[Y(D((0,0)))|E=1,T=1]-E[Y(D((0,0)))|E=1,T=0]$
$-\{E[Y(D((0,0)))|E=0,T=1]-E[Y(D((0,0)))|E=0,T=0]\Bigg{\}}$
$=$ $E[Y(D((0,1)))-Y(D((0,0)))|E=1,T=1]$
$=$ $E[(D((0,1))-D((0,0)))\cdot(Y(1)-Y(0))|E=1,T=1]$
$=$ $E[Y(1)-Y(0)|T=1,E=1,CM^{Z}]\cdot Pr(CM^{Z}|E=1,T=1).$
$=$ $LATET\cdot Pr(CM^{Z}|E=1,T=1).$ (15)
The second equality follows from the simple manipulation. The third equality follows from Assumption 16. The fourth equality follows from Assumption 18 and the simple calculation. The fifth equality follows from the Law of Iterated Expectations and Assumption 15.
Combining the result (14) with (15), we obtain
$w_{DID}$ $=\frac{LATET\cdot Pr(CM^{Z}|E=1,T=1)}{Pr(CM^{Z}|E=1,T=1)}$
$=LATET.$
B.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 7 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are independent and identically distributed (i.i.d).
In multiple time period and multiple cohort settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
@@ B.2.2 Multiple time periods
When repeated cross section data are available and we have multiple time periods, the staggered DID-IV designs consist of Assumptions 7, 10, 11 (in Section 3), Assumption 14 (Exclusion restriction), and the following parallel trends assumptions.
@@
[…]
@@
[…]
B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
We discuss the DID-IV identification strategy in two periods with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. We consider two types of the no anticipation assumption on potential outcomes $Y_{i,t}(d,z)$ instead of imposing Assumption 1 (No carryover assumption). We then define the expected time gain and formalize the parallel trends assumption in the outcome as in Assumption 6 in this setting. Finally, we show that a Wald-DID estimand identifies the LATET’ (see Section 4) under suitable assumptions similar to the ones in Subsection 2.3. We consider the same setting as in Subsection 2.1 except that we do not adopt Assumption 1.
First, we make the following no anticipation assumptions on the outcome.
@@ B.3   Introducing the treatment path in $2\times 2$ DID-IV designs
[…]
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$. Note that this assumption corresponds to the no anticipation assumption usually made in the common DID set up as in Roth et al. (2023) (see Table 3 and Table 4 in Subsection 2.5) [^9] It is inconsequential whether the instrument path $z$ is included or not in potential outcomes because we have $d=z$ in the common DID set up..
@@
[…]
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$. Assumption 22 is inherent to the DID-IV identification strategy when we introduce the treatment path.
We can rationalize Assumption 21 and 22 by noting that both of them and the no anticipation assumption proposed by the recent DID literature restrict the anticipatory behavior of the units belonging to the exposed group and the compliers $CM^{Z}$.
We now make the parallel trends assumption in the outcome suitable for the set up considered here. We first define the similar notation as in Subsection 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
@@
$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},$
$\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
@@
As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the homogeneous expected time gain between exposed and unexposed groups. The following assumption makes such a restriction in the current set up.
@@
The expected time gain between exposed and unexposed groups is the same; that is, we have $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
The theorem below shows that a Wald-DID estimand is equal to the target parameter $LATET^{\prime}$ under Assumption 2-5 and Assumption 21-23.
@@
Suppose Assumption 2-5 and Assumption 21-23 holds. Then, a Wald-DID estimand corresponds to $LATET^{\prime}$; that is, we have
$\frac{E[Y_{1}-Y_{0}|E=1]-E[Y_{1}-Y_{0}|E=0]}{E[D_{1}-D_{0}|E=1]-E[D_{1}-D_{0}|E=0]}=LATET^{\prime}.$
@@
The second equality holds from the fact that we can partition the compliers $CM^{Z}$ in an exposed group into $(CM^{Z}\land DF^{T})$ and $(CM^{Z}\land NT^{T})$ (see Table 1).
Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
@@
Next, we decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
$E[Y_{1}-Y_{0}|E=0]$
$=E[Y_{1}((1,1))-Y_{0}((1,1))|E=0,AT^{T}]Pr(AT^{T}|E=0)$
$+E[Y_{1}((0,1))-Y_{0}((0,1))|E=0,CM^{T}]Pr(CM^{T}|E=0)$
$+E[Y_{1}((0,0))-Y_{0}((0,0))|E=0,NT^{T}]Pr(NT^{T}|E=0)$
$+E[Y_{1}((1,0))-Y_{0}((1,0))|E=0,DF^{T}]Pr(DF^{T}|E=0)$
$=\bar{\Delta}_{0}.$ (18)
The first equality follows from Table 2 and the second follows from the definition of the expected time gain $\bar{\Delta}_{0}$.
@@
Appendix C Proofs in Section 4
C.1   Proof of theorem 4
First, we prove the consistency of our SLS estimator. $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ take the following form:
$\equiv\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}.$
$\equiv\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}.$
We assume that $\hat{\alpha}^{e,l}$ and $\hat{\pi}^{e,l}$ are well defined. More precisely, we assume that cohort $e$ is non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}=e\}>0$ and there exist some sets of the unexposed cohort $U\subseteq\{u:e-1<u\leq T\}$ that are non-empty; that is, $\sum_{i=1}^{n}\mathbf{1}\{E_{i}\in U\}>0$. This assumption guarantees that each Wald-DID estimator
$\widehat{CLATT}_{e,e+l}$ $=\frac{\hat{\alpha}^{e,l}}{\hat{\pi}^{e,l}}$
$=\frac{\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}}{\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}}.$
is also well defined.
First we derive the probability limit of $\widehat{CLATT}_{e,e+l}$. By the Law of Large Numbers and Slutsky’s theorem, we have
@@ C.1   Proof of theorem 4
$=E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e],$
where the second equality follows from the Law of Iterated Expectations.
By similar argument, we have
[…]
From Slutsky’s theorem, we have
[…]
Lastly, we show that the right hand side of (20) is equal to $CLATT_{e,e+l}$.
For the numerator, we have
$E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}=e]-\sum_{e\in U}E[Y_{i,e+l}(D_{i,e+l}^{\infty})-Y_{i,e-1}(D_{i,e-1}^{\infty})|E_{i}\in U]Pr(E_{i}=e|E_{i}\in U)$
$=$ $E[Y_{i,e+l}(D_{i,e+l}^{e})-Y_{i,e+l}(D_{i,e+l}^{\infty})|E_{i}=e]$
The third equality follows from Assumption 8, 9 and 11. The fourth equality follows from Assumption 13 and the fifth follows from Assumption 10.
For the denominator, by similar calculations,
$E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]$
$=E[D_{i,e+l}^{e}-D_{i,e+l}^{\infty}|E_{i}=e]+E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]-\sum_{e\in U}E[D_{i,e+l}^{\infty}-D_{i,e-1}^{\infty}|E_{i}=e]Pr(E_{i}=e|E_{i}\in U)$
$=Pr(CM_{e,e+l}|E_{i}=e),$ (22)
where the final equality follows from Assumption 10 and Assumption 12.
Combining (21) with (22), we obtain the desirable result:
$\frac{E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}\in U]}{E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]-E[D_{i,e+l}-D_{i,e-1}|E_{i}\in U]}$
$=\frac{E[Y_{i,e+l}(1)-Y_{i,e+l}(0)|E_{i}=e,CM_{e,e+l}]Pr(CM_{e,e+l}|E_{i}=e)}{Pr(CM_{e,e+l}|E_{i}=e)}$
$=CLATT_{e,e+l}.$
Next, we prove that our STS estimator is asymptotically normal, deriving its influence function.
We define $\theta_{e,l}$ to be
$\equiv\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}.$
The following fact, also found in de Chaisemartin and D’Haultfœuille (2018), is repeatedly used in the derivation.
Fact 1.
If
we have
First we derive the influence function of $\sqrt{n}(\hat{\alpha}_{e,l}^{1}+\hat{\alpha}_{e,l}^{2}-(\alpha_{e,l}^{1}+\alpha_{e,l}^{2}))$. We note that
$\sqrt{n}(E_{N}[\mathbf{1}\{E_{i}=e\}]-E[\mathbf{1}\{E_{i}=e\}])=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\mathbf{1}\{E_{i}=e\}-E[\mathbf{1}\{E_{i}=e\}]).$
and
$\sqrt{n}(E_{N}[(Y_{i,e+l}-Y_{i,e-1})\mathbf{1}\{E_{i}=e\}]-E[(Y_{i,e+l}-Y_{i,e-1})\mathbf{1}\{E_{i}=e\}])$
$=\frac{1}{\sqrt{n}}\sum_{i}^{n}((Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}-E[(Y_{i,e+l}-Y_{i,e-1})\cdot\mathbf{1}\{E_{i}=e\}]).$
From the observations and using Fact 1, we obtain
$\sqrt{n}(\hat{\alpha}_{e,l}^{1}-\alpha_{e,l}^{1})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\frac{\mathbf{1}\{E_{i}=e\}\cdot\left[(Y_{i,e+l}-Y_{i,e-1})-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]}+o_{p}(1).$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}\zeta_{i,e,l}+o_{p}(1).$
By symmetry, we also obtain the influence function of $\sqrt{n}(\hat{\alpha}_{e,l}^{2}-\alpha_{e,l}^{2})$:
$\sqrt{n}(\hat{\alpha}_{e,l}^{2}-\alpha_{e,l}^{2})$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}\frac{\mathbf{1}\{E_{i}=U\}\cdot\left[(Y_{i,e+l}-Y_{i,e-1})-E[Y_{i,e+l}-Y_{i,e-1}|E_{i}=U]\right]}{E[\mathbf{1}\{E_{i}=U\}]}+o_{p}(1).$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}\phi_{i,e,l}+o_{p}(1).$
This implies that
$\sqrt{n}(\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}-(\alpha_{e,l}^{1}-\alpha_{e,l}^{2}))$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\zeta_{i,e,l}-\phi_{i,e,l})+o_{p}(1)$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}a_{i,e,l}+o_{p}(1).$ (23)
Next we derive the influence function of $\sqrt{n}(\hat{\pi}_{e,l}^{1}+\hat{\pi}_{e,l}^{2}-(\pi_{e,l}^{1}+\pi_{e,l}^{2}))$. From the symmetry, by replacing $Y$ with $D$, we obtain the following:
$\sqrt{n}(\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}-(\pi_{e,l}^{1}-\pi_{e,l}^{2}))$ $=\frac{1}{\sqrt{n}}\sum_{i}^{n}(\eta_{i,e,l}-\xi_{i,e,l})+o_{p}(1)$
$\equiv\frac{1}{\sqrt{n}}\sum_{i}^{n}b_{i,e,l}+o_{p}(1),$ (24)
where we define
$\eta_{i,e,l}\equiv\frac{\mathbf{1}\{E_{i}=e\}\cdot\left[(D_{i,e+l}-D_{i,e-1})-E[D_{i,e+l}-D_{i,e-1}|E_{i}=e]\right]}{E[\mathbf{1}\{E_{i}=e\}]},$
$\xi_{i,e,l}\equiv\frac{\mathbf{1}\{E_{i}=U\}\cdot\left[(D_{i,e+l}-D_{i,e-1})-E[D_{i,e+l}-D_{i,e-1}|E_{i}=U]\right]}{E[\mathbf{1}\{E_{i}=U\}]}.$
Using (23), (24) and Fact 1, by some calculations, we obtain
$\sqrt{n}(\widehat{CLATT}_{e,e+l}-CLATT_{e,e+l})$ $=\sqrt{n}(\frac{\hat{\alpha}_{e,l}^{1}-\hat{\alpha}_{e,l}^{2}}{\hat{\pi}_{e,l}^{1}-\hat{\pi}_{e,l}^{2}}-\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}})$
$=\frac{1}{\sqrt{n}}\sum_{i}^{n}\psi_{i,e,l}+o_{p}(1),$
where the influence function $\psi_{i,e,l}$ is:
$\psi_{i,e,l}$
$=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}\left(\zeta_{i,e,l}-\phi_{i,e,l}-\frac{\alpha_{e,l}^{1}-\alpha_{e,l}^{2}}{\pi_{e,l}^{1}-\pi_{e,l}^{2}}(\eta_{i,e,l}-\xi_{i,e,l})\right)$
Therefore, the asymptotic variance $V(\psi_{i,e,l})$ is
$V(\psi_{i,e,l})$
$=\frac{1}{(\pi_{e,l}^{1}-\pi_{e,l}^{2})^{2}}\Bigg{(}\frac{E\big{[}\mathbf{1}\{E_{i}=e\}\left[\delta_{i,e,l}^{p}-E[\delta_{i,e,l}^{p}|E_{i}=e]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=e\}]^{2}}+\frac{E\big{[}\mathbf{1}\{E_{i}=U\}\left[\delta_{i,e,l}^{p}-(E[\delta_{i,e,l}^{p}|E_{i}=U]\right]^{2}\big{]}}{E[\mathbf{1}\{E_{i}=U\}]^{2}}\Bigg{)}.$
C.2   Repeated cross sections
We also present the influence function of our STS estimator when repeated cross section data are available. [^10] The derivation here is essentially the same in de Chaisemartin and D’Haultfœuille (2018), who presented the influence function of the Wald-DID estimator in repeated cross section settings with two periods and two groups.. Let $\theta_{e,l}^{r}$ define
@@ C.2   Repeated cross sections
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Using similar arguments in the proof of Theorem 4, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 7, 10, 11, 14, 19, and 20, and have the following influence function:
$\psi_{i,e,l}^{r}$ $=\frac{1}{\pi_{e,l}^{1}-\pi_{e,l}^{2}-[\gamma_{e,l}^{1}-\gamma_{e,l}^{2}]}$
$\times$ $\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$

### v5

@@ A.1   Non-binary, ordered treatment
For the numerator of the Wald DID estimand, we have
@@
This follows from Assumption 4 and Assumption 6.
@@
For the denominator of the Wald-DID estimand, by similar argument, we obtain:
@@
This follows from Assumption 4 and 5.
@@
A.2   Repeated cross sections
Here, we present our identification results under repeated cross section settings. First, we consider two groups and two periods settings. Next, we consider multiple period settings with staggered instrument.
A.2.1 Two time periods
Let $Y_{i}$ and $D_{i}$ denote the outcome and the treatment for unit $i$. Let $Z_{i}$ denote the instrument path for unit $i$: $Z_{i}=(0,0)$ if unit $i$ is not exposed to the instrument and $Z_{i}=(0,0)$ if unit $i$ is exposed to the instrument. Let $E_{i}\in\{0,1\}$ denote the group indicator for unit $i$: $E_{i}=1$ if $Z_{i}=(0,1)$ and $E_{i}=0$ if $Z_{i}=(0,0)$. Let $T_{i}\in\{0,1\}$ denote the binary indicator for time. For all $z\in\mathcal{S}(Z)$, let $Y_{i}(0,z),Y_{i}(1,z)$, and $D_{i}(z)$ denote potential outcomes and potential treatment choices for unit $i$. Let $D_{i,t}(z)$ denote potential treatment choices under $T_{i}=t$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In two periods and two groups settings, our target parameter is the local average treatment effect on the treated in period $1$:
@@ A.2.1 Two time periods
Assumption 19 (Relevance condition).
$E[D_{i}|E_{i}=1,T_{i}=1]-E[D_{i}|E_{i}=1,T_{i}=0]$
-- $(E[D_{i}|E_{i}=0,T_{i}=1]-E[D_{i}|E_{i}=0,T_{i}=0])>0.$
By similar arguments in the proof of Theorem 1 (Theorem 4), one can show that the Wald-DID estimand is equal to the LATET (ACRT) in period $1$ under Assumptions 14-19 (thus, we omit the proof for brevity).
A.2.2 Multiple time periods
Let $T_{i}\in\{1,\dots,T\}$ denote the time period when unit $i$ is observed and let $Z_{i,t}$ denote the instrument status under $T_{i}=t$. We adopt Assumption 8 in Section 3; that is, we assume the staggered assignment of the instrument across units. In addition, we adopt Assumption 14 (Exclusion restriction), which allows us to write $Y_{i}(D_{i},z)=Y_{i}(D_{i})$. Let $E_{i}\in\{1,\dots,T,\infty\}$ denote the cohort which unit $i$ belongs to. Let $Y_{i}(D_{i}^{e})$ and $D_{i}^{e}$ denote potential outcomes and potential treatment choices when unit $i$ belongs to cohort $e$. Let $D_{i,t}^{e}$ denote potential treatment choices under $T_{i}=t$ when unit $i$ belongs to cohort $e$. We assume that $\{Y_{i},D_{i},Z_{i},E_{i},T_{i}\}_{i=1}^{n}$ are i.i.d.
In multiple periods and cohorts settings, our target parameter is the cohort specific local average treatment effect on the treated in period $T=e+l$:
@@ A.2.2 Multiple time periods
When repeated cross section data are available, the staggered DID-IV designs consist of Assumptions 8, 11, 12 (in Section 3), Assumption 14, and the following parallel trends assumptions and relevance condition.
@@
$\text{For all}\hskip 2.84526pts<t,\hskip 2.84526ptE[D_{i}^{\infty}|E_{i}=e,T_{i}=t]-E[D_{i}^{\infty}|E_{i}=e,T_{i}=s]\hskip 2.84526pt\text{is identical for all}\hskip 2.84526pte\in\mathcal{S}(E_{i}).$
@@
For $\text{For all}\hskip 5.69054pts<t,\hskip 5.69054ptE[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=t]-E[Y_{i}(D_{i}^{\infty})|E_{i}=e,T_{i}=s]$
is identical for $\text{is identical for all}\hskip 5.69054pte\in\mathcal{S}(E_{i}).$
Assumption 22 (Relevance condition in multiple time periods).
$\forall e\in\mathcal{S}(E_{i}),\hskip 5.69054pt\forall l\geq 0,\hskip 5.69054ptE[D_{i}-D^{\infty}_{i}|E_{i}=e,T_{i}=e+l]>0.$
A.3   Introducing the treatment path in $2\times 2$ DID-IV designs
We discuss the DID-IV identification strategy in two periods and two groups with a binary treatment, introducing the treatment path in potential outcomes $Y_{i,t}(d,z)$. In this setting, our target parameter is the LATET’ defined below.
Definition.
$LATET^{\prime}\equiv w_{0}\cdot\Delta_{0}+w_{1}\cdot\Delta_{1}$
where $\Delta_{0}$ and $\Delta_{1}$ are:
The weights $w_{0}$ and $w_{1}$ are:
We consider the identifying assumptions for the Wald-DID estimand to identify the LATET’. First, we make the following no anticipation assumptions on the outcome instead of imposing Assumption 1 (No carryover assumption).
@@ A.3   Introducing the treatment path in $2\times 2$ DID-IV designs
$\forall z\in\mathcal{S}(Z),Y_{i,0}((0,1),z)=Y_{i,0}((0,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land NT^{T}).$
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((0,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land NT^{T}$.
@@
$\forall z\in\mathcal{S}(Z),Y_{i,0}((1,1),z)=Y_{i,0}((1,0),z)\hskip 5.69054pt\text{for all}\hskip 5.69054pti\hskip 5.69054pt\text{with}\hskip 5.69054ptE_{i}=1\hskip 5.69054pt\text{and}\hskip 5.69054pt(CM^{Z}\land DF^{T}).$
This assumption requires that the potential outcome at period 00 is the same as the baseline outcome $Y_{i,0}((1,0),z)$, given the instrument path $z$, for those assigned to the instrument $z=(0,1)$ and belonging to the type $CM^{Z}\land DF^{T}$.
We can rationalize Assumptions 23-24 by considering the similarity between these assumptions and the no anticipation assumption made in the recent DID literature; all of them restrict the anticipatory behavior for those who belong to an exposed group (treatment group) and are affected by a policy shock (the treatment in DID designs).
Next, we make the parallel trends assumption in the outcome suitable for the set up considered here. We first define the similar notation as in Subsection 2.4. Let $\bar{\Delta}_{g^{t},e}$ denote the expectation of the time trends in each groups in group $E_{i}=e$:
@@
$\bar{\Delta}_{0}\equiv\sum_{g^{t}}w_{g^{t},0}\bar{\Delta}_{g^{t},0},\hskip 5.69054pt\bar{\Delta}_{1}\equiv\sum_{g^{t}}w_{g^{t},1}\bar{\Delta}_{g^{t},1},$
@@
As noted in Subsection 2.4, the parallel trends assumption in the outcome can be interpreted as requiring the same expected time gain between exposed and unexposed groups. The following assumption makes such a restriction in the current set up.
@@
The expected time gain between exposed and unexposed groups is the same: $\bar{\Delta}_{0}=\bar{\Delta}_{1}$.
The theorem below shows that the Wald-DID estimand is equal to the LATET’ under Assumptions 2-5, 7 and Assumptions 23-25.
@@
Suppose Assumptions 2-5, 7 and Assumptions 23-25 hold. Then, the Wald-DID estimand corresponds to the LATET’.
@@
Next, for the numerator, we first decompose $E[Y_{1}-Y_{0}|E=0]$ as follows.
$E[Y_{1}-Y_{0}|E=0]=\bar{\Delta}_{0}.$ (16)
This follows from Table 2 and the definition of the expected time gain $\bar{\Delta}_{0}$.
Finally, we decompose $E[Y_{1}-Y_{0}|E=1]$ as follows.
@@
Appendix B Deriving the influence function in repeated cross section settings
@@
In repeated cross section settings, our STS estimator is the sample analog of $\theta_{e,l}^{r}$, which we denote $\hat{\theta}_{e,l}^{r}$. Using similar arguments in the proof of Theorem 3, one can show that $\hat{\theta}_{e,l}^{r}$ is consistent for $CLATT_{e,e+l}$ under Assumptions 8, 11, 12, 14, 20-22, and have the following influence function:
$\psi_{i,e,l}^{r}=\frac{1}{\pi_{e,e+l}-\pi_{e,e-1}-[\gamma_{U,e+l}-\gamma_{U,e-1}]}$
×$\times\Bigg{(}\frac{\mathbf{1}_{e,e+l}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e+l]\right]}{E[\mathbf{1}_{e,e+l}]}-\frac{\mathbf{1}_{e,e-1}\cdot\left[\delta_{i,e,l}-E[\delta_{i,e,l}|E_{i}=e,T_{i}=e-1]\right]}{E[\mathbf{1}_{e,e-1}]}$